# Optimizing wheel loader performance — an end-to-end approach

### v2

@@ 4.4.3 Nominal strategy
For comparison, we also test the strategy of always selecting the dig location that has the lowest transportation cost, irrespective of the loading performance. This is accomplished with $Q^{d=1}_{n}=Q_{\text{nominal}}\equiv\bm{w}^{\mathsf{T}}\left(\bm{\mathcal{P}_{n}}^{\text{V1}}+\bm{\mathcal{P}_{n}}^{\text{V2}}\right)$. We refer to this as the nominal strategy as it is a standard method [24] and the natural choice when there is no access to a world model.
@@ 4.5 Computational time
The computational time is what limits us from computing the optimal solution by exhaustive search. We register the computational time for analyzing a single loading cycle. It includes predicting the new pile state $\bm{H}^{\prime}$ and total performance $\bm{\mathcal{P}}$, which involves updating $\bm{a}^{\text{load}}$ accorging to Eq. (9) (dominant part) and computing and integrating the V-turns given a selected dig location $\bm{x}^{\text{dig}}$ as described in Sec. 4.2. Table 1 shows the average computational time from profiling the loading predictions made on a desktop computer with an Intel(R) Core(TM) i7-8700K, 3.70 GHz, 32 GB RAM on a Windows 64-bit system and NVIDIA GeForce RTX 2070 SUPER.
@@ 5 Results
We test the look-ahead search algorithm considering a wheel loader performing a task of $N=15$ sequential loading cycles. First, we analyze the greedy strategy (prediction horizon $d=1$) and compare it using the load optimizing and the nominal strategies. Next, we analyze the look-ahead tree search algorithm with different search depths to examine the effect of the prediction horizon.
The initial pile surface $\bm{H}_{1}$ features a trapezoidal prism, 1.8 m tall with a $30^{\circ}$ slope at the front, with Perlin noise [25] added to the surface. A representative initial pile surface is illustrated in Fig. 6.
@@ 5.1 Greedy strategy
The greedy strategy is 8 % more productive and energy efficient than the nominal strategy, using nearly the same time to complete 15 cycles. The maximum loading strategy is only 1 % more productive than the greedy strategy in loading tasks but spends around 8 % more time and energy during V-turns. The analysis also demonstrates that the evaluation function works effectively.
@@ 6 Discussion
With the look-ahead search method, it is possible to find action sequences that maximize the total performance over long horizons by increasing the pile state quality for future loadings. This method adapts to the truck’s location and can do so even if it would vary provided that the location is known in advance. In our experiments, the optimal action sequence can be (before any serious code optimization) computed in 43 s with a look-ahead search with depth $d=4$, which is close to the average loading cycle time. Future work should consider other soil properties, such as cohesive or inhomogeneous soil. The resulting pile state is likely to depend more strongly on the selected loading actions than it does for gravel, for which the pile always tends to settle with a local slope near the natural angle of repose. Capturing this with a world model would be more demanding and might benefit from a longer horizon than $d=4$. That would motivate substantial improvements in the computational efficiency of the look-ahead search.

### v3

@@ 4.4.3 Nominal strategy
For comparison, we also test the strategy of always selecting the dig location that has the lowest transportation cost, irrespective of the loading performance. This is accomplished with $Q^{d=1}_{n}=Q_{\text{nominal}}\equiv\bm{w}^{\mathsf{T}}\left(\bm{\mathcal{P}_{n}}^{\text{V1}}+\bm{\mathcal{P}_{n}}^{\text{V2}}\right)$. We refer to this as the nominal strategy as it is a standard method [hong2018path] and the natural choice when there is no access to a world model.
@@ 4.5 Computational time
The computational time is what limits us from computing the optimal solution by exhaustive search and instead approximate the evaluation function by Eq. (4.4) at higher search depths. We register the computational time for analyzing a single loading cycle. It includes predicting the new pile state $\bm{H}^{\prime}$ and total performance $\bm{\mathcal{P}}$, which involves updating $\bm{a}^{\text{load}}$ accorging to Eq. (9) (dominant part) and computing and integrating the V-turns given a selected dig location $\bm{x}^{\text{dig}}$ as described in Sec. 4.2. Table 1 shows the average computational time from profiling the loading predictions made on a desktop computer with an Intel(R) Core(TM) i7-8700K, 3.70 GHz, 32 GB RAM on a Windows 64-bit system and NVIDIA GeForce RTX 2070 SUPER.
@@ 5 Results
We test the look-ahead search algorithm considering a wheel loader performing a task of $N=15$ sequential loading cycles. First, we analyze the greedy strategy (prediction horizon $d=1$) and compare it using the load optimizing and the nominal strategies. Next, we analyze the look-ahead tree search algorithm with different search depths to examine the effect of the prediction horizon. The weights were fixed at $\bm{w}=[2,1,1]$ such that production (loaded mass) and cost (loading time, mechanical work) contribute equally to the overall loading performance.
The initial pile surface $\bm{H}_{1}$ features a trapezoidal prism, 1.8 m tall with a $30^{\circ}$ slope at the front, with Perlin noise [perlin1985image] added to the surface to avoid ending up in the same local minima. A representative initial pile surface is illustrated in Fig. 6.
@@ 5.1 Greedy strategy
The greedy strategy is 8 % more productive and energy efficient than the nominal strategy, using nearly the same time to complete 15 cycles. The maximum loading strategy is 1 % more productive than the greedy strategy in loading tasks but spends around 8 % more time and energy during V-turns. That is an expected result, as transportation time and energy do not contribute to the evaluation function of the maximum loading strategy. The analysis demonstrates that the evaluation function works effectively.
@@ 6 Discussion
With the look-ahead search method, it is possible to find action sequences that maximize the total performance over long horizons by increasing the pile state quality for future loadings. This method adapts to the truck’s location and can do so even if it would vary provided that the location is known in advance. In our experiments, the optimal action sequence can be (before any serious code optimization) computed in 43 s with a look-ahead search with depth $d=4$, which is close to the average loading cycle time. Future work should consider other soil properties, such as cohesive or inhomogeneous soil. The resulting pile state is likely to depend more strongly on the selected loading actions than it does for gravel, for which the pile always tends to settle with a local slope near the natural angle of repose. To support more realistic conditions, one should extend the framework to support non-flat ground and spillage. Capturing this with a world model would be more demanding and might benefit from a longer horizon than $d=4$. That would motivate substantial improvements in the computational efficiency of the look-ahead search. In this study, we employed a set of specific weight factors that equally balance productivity (mass per unit time) and energy efficiency (mass per unit energy). It would be interesting, in future work, to explore the impact of different weight factors on the result, e.g., what digging behaviour emerges when energy consumption is heavily penalized.